# Chart-R1: Chain-of-Thought Supervision and Reinforcement for Advanced Chart Reasoner

### v1

@@ 1 Introduction
Figure 1: Comparison of existing VLM performance on Chart understanding and reasoning benchmarks. In the complex chart reasoning task, Qwen2.5-VL-7B generate wrong thinking process, whereas Chart-R1 thinks and answers correctly.
@@ 2.2 Long Reasoning VLMs
Recently, with the success of DeepSeek-R1 Guo et al. (2025), many works have attempted to enhance the LLM reasoning ability via rule-based reward and RL Shao et al. (2024). In the vision-language domain, recent works follow the DeepSeek-R1 method to enhance the long-chain reasoning capacity of VLMs Shen et al. (2025); Wang et al. (2025). For example, Vision-R1 Huang et al. (2025) and R1-OneVision Yang et al. (2025) apply Group Relative Policy Optimization (GRPO) with multimodal reasoning data to enable VLMs for long reasoning. MMEureka Meng et al. (2025) and R1-Zero Liu et al. (2025) further advance the visual long-term reasoning with improved RL training strategies. Point-RFT Ni et al. (2025) utilizes grounded CoT reasoning for visual understanding, but it just utilize ChartQA for RL which limits the final model reasoning capacity.
@@ 2.3 Chart Understanding and Reasoning
A variety of training datasets and evaluation benchmarks have been developed to improve VLM performance on chart-related tasks Xia et al. (2024); Shi et al. (2024); Zhao et al. (2025a); Wu et al. (2025). Previous works generally focus on description tasks, for example, ChartQA Masry et al. (2022), PlotQA Methani et al. (2020) and Chart-to-text Kantharaj et al. (2022) mainly train and evaluate the capacities of the models on extracting information from the chart. While numerous relevant works exist, the challenge in the description tasks is predominantly driven by chart complexity. Recent works such as Charxiv Wang et al. (2024) and CharMuseum Tang et al. (2025) introduce more challenging reasoning tasks, demanding that models think before answering. Unlike descriptive tasks, reasoning tasks present a dual challenge, originating from both the perceptual complexity of the chart and the reasoning depth required by questions.
@@ 4.1 Implementation Details
InternVL3-8B 86.6 37.6 - -
InternVL3-38B 89.2 46.4 - -
@@ 4.2 Experiment Settings
Baselines We compare our proposed Chart-R1 with existing models in three setups: (1) Proprietary models include GPT-4oOpenAI (2024), Gemini-1.5-(Flash, Pro)Team et al. (2023), Gemini-2.5-Flash and Claude-3.5-SonnetAnthropic (2024). (2) General-domain open-source VLMs including InternVL3(8B, 38B) Zhu et al. (2025), Qwen2.5-VL(7B) Bai et al. (2025). (3) Chart-domain VLMs including ChartLlama Han et al. (2023), TinyChart Zhang et al. (2024), ChartGemma Masry et al. (2024b) and ChartResoner Jia et al. (2025).
@@ 4.4 Ablation Study
Visualization We conclude with qualitative case studies where our Chart-R1 model successfully generates detailed reasoning and correct answers for complex questions. In these same instances, the baseline Qwen2.5VL-7B model fails, directly demonstrating the superior performance and more advanced reasoning capabilities of our approach.
Figure 6: The visualization result of a case study that Chart-R1 (w/ Chart-COT) answer correctly, but Qwen2.5VL-7B and Chart-R1 (w/o Chart-COT) fails.

### v2

@@ 1 Introduction
Figure 1: Comparison of existing VLM performance on chart understanding and reasoning benchmarks. In the complex chart reasoning task, Qwen2.5-VL-7B generates a wrong thinking process, whereas Chart-R1 thinks and answers correctly.
@@ 2.2 Long Reasoning VLMs
Recently, with the success of DeepSeek-R1 Guo et al. (2025), many works have attempted to enhance the LLM reasoning ability via rule-based reward and RL Shao et al. (2024). In the vision-language domain, recent works follow the DeepSeek-R1 method to enhance the long-chain reasoning capacity of VLMs Shen et al. (2025); Wang et al. (2025); Qiu et al. (2025). For example, Vision-R1 Huang et al. (2025) and R1-OneVision Yang et al. (2025) apply Group Relative Policy Optimization (GRPO) with multimodal reasoning data to enable VLMs for long reasoning. MMEureka Meng et al. (2025b) and R1-Zero Liu et al. (2025) further advance the visual long-term reasoning with improved RL training strategies. Point-RFT Ni et al. (2025) utilizes grounded CoT reasoning for visual understanding, but it just utilize ChartQA for RL which limits the final model reasoning capacity.
@@ 2.3 Chart Understanding and Reasoning
A variety of training datasets and evaluation benchmarks have been developed to improve VLM performance on chart-related tasks Xia et al. (2024); Shi et al. (2024); He et al. (2024); Zhao et al. (2025a); Wu et al. (2025). Previous works generally focus on description tasks, for example, ChartQA Masry et al. (2022), PlotQA Methani et al. (2020) and Chart-to-text Kantharaj et al. (2022) mainly train and evaluate the capacities of the models on extracting information from the chart. While numerous relevant works exist, the challenge in the description tasks is predominantly driven by chart complexity. Recent works such as Charxiv Wang et al. (2024) and CharMuseum Tang et al. (2025) introduce more challenging reasoning tasks, demanding that models think before answering. Unlike descriptive tasks, reasoning tasks present a dual challenge, originating from both the perceptual complexity of the chart and the reasoning depth required by questions.
@@ 4.1 Implementation Details
Phi-3.5-Vision 81.8 32.7 24.73 31.08 / 24.32
DeepSeek-VL2 86.0 - 16.28 23.15 / 20.29
InternVL3-8B 86.6 37.6 - 37.51 / 31.73
InternVL3-38B 89.2 46.4 - 46.09 / 38.36
@@ 4.2 Experiment Settings
Baselines We compare our proposed Chart-R1 with existing models in three setups: (1) Proprietary models include GPT-4oOpenAI (2024), Gemini-1.5-(Flash, Pro)Team et al. (2023), Gemini-2.5-Flash and Claude-3.5-SonnetAnthropic (2024). (2) General-domain open-source VLMs including Phi 3.5-Vision Abdin et al. (2024), DeepSeek-VL2 Wu et al. (2024), InternVL3(8B, 38B) Zhu et al. (2025) and Qwen2.5-VL(7B) Bai et al. (2025). (3) Chart-domain VLMs including ChartLlama Han et al. (2023), TinyChart Zhang et al. (2024), ChartGemma Masry et al. (2024b) and ChartResoner Jia et al. (2025).
@@ 4.4 Ablation Study
Visualization We conclude with qualitative case studies where our Chart-R1 model successfully generates detailed reasoning and correct answers for complex questions in Figures 5 and 6. In these same instances, the baseline Qwen2.5VL-7B model fails, directly demonstrating the superior performance and more advanced reasoning capabilities of our approach. Without training on the Chart-COT stage, Chart-R1 w/o Chart-COT also fails to answer problems in Figure 6, which shows the importance of our proposed two-stage training.
Figure 6: The visualization result of a case study that Chart-R1 (w/ Chart-COT) answers correctly, but Qwen2.5VL-7B and Chart-R1 (w/o Chart-COT) fail.
@@ 5 Conclusion
Appendix A More Training Details
In this section, we provide more details about Chart-R1’s training process, including Chart-COT and Chart-RFT.
Chart-COT We use Qwen2.5VL-7B-Instruct as the initial model and perform supervised fine-tuning using LLaMA-Factory Zheng et al. (2024). We train the model on the 228k ChartRQA-SFT dataset for one epoch. During training, we freeze the vision tower and multi-modal projector parameters and tune the LLM. The learning rate is set to 1e-5, with a warm-up ratio of 0.1 and batch size of 48. The training process costs 3 hours on 24 H800 GPUs.
Chart-RFT For the RFT stage, we use the fine-tuned model from the Chart-COT stage. We adopt the MM-EUREKA Meng et al. (2025a) framework based on OpenRLHF for training. The model is trained for 3 episodes using 30k ChartQA and 30k ChartRQA-RL. We set the rollout batch size and the training batch size to 128, with each sample generating 8 rollouts. The temperature for model generation is set to 1, and we exclude KL divergence in the loss calculation. The learning rate is set to 1e-6, with a warm-up ratio of 0.03, while freezing the vision tower during training. Following the default setting for instruction models, the format reward coefficient is set to 0.5. We employ the online filtering strategy with lower and upper bounds of 0.1 and 0.9, respectively. The training process costs 30 hours on 24 H800 GPUs.
Appendix B Benchmark Details
ChartQA Masry et al. (2022) focuses on chart question answering with complex reasoning questions that involve logical and arithmetic operations. Following the settings in the original paper, we evaluate models on the test set reporting overall accuracy scores across both human-written (ChartQA-H) and machine-generated (ChartQA-M) question subsets.
CharXiv Wang et al. (2024) presents a comprehensive evaluation suite with natural, challenging, and diverse charts from arXiv papers to provide a more realistic assessment of chart understanding capabilities. We evaluate models on the Reasoning Questions (CharXiv-RQ) subset, which requires synthesizing information across complex visual elements in charts. Following the original paper, we use GPT-assisted evaluation to assess model responses.
ChartQAPro Masry et al. (2025) introduces a diverse benchmark with various chart types, including infographics and dashboards, and question formats that better reflect real-world challenges. We evaluate models using Chain-of-Thought (CoT) prompting in the original paper and report overall accuracy across five question types.
Appendix C ChartRQA Analysis
We count the quantity and distribution of different chart types across the training and test sets of ChartRQA, as detailed in Table A. The distribution among the various types to be well-balanced. Furthermore, Figures A and  B provide visualization examples of 24 chart types from the ChartRQA dataset, showcasing both single-chart and multi-chart formats, respectively.
Table A: The detailed chart types and corresponding quantities in our proposed ChartRQA train and test set. ChartRQA contains 24 chart types, each of which contains approximate samples.
Split Bar Line ErrorBar Heatmap Box Scatter Histogram Radar 3D
Train 11,850 10,752 11,838 8,993 12,112 10,299 15,856 9,483 9,746
Test 100 88 83 60 103 76 116 46 65
Split Pie ErrorPoint Violin Area Bubble Multi-axes Ring Rose Treemap
Train 17,812 10,814 12,571 9,175 8,996 10,776 12,726 10,533 9,850
Test 103 68 116 75 51 61 54 61 64
Split Bar_num Contour Density Graph Quiver Funnel Total
Train 12,150 10,291 12,860 8,764 9,955 227 258,429
Test 64 67 77 47 52 5 1,702
Figure A: Single-chart samples of 24 chart types from ChartRQA.
Figure B: Multi-chart samples of 24 chart types from ChartRQA.
Appendix D Prompts
To enhance transparency and reproducibility, we provide the exact prompts used for dataset generation and evaluation.
Figure C illustrates the prompt used for code generation. We utilize real table data as input, select one chart type from the 24 predefined chart types, and sample a code example corresponding to that chart type to generate the plotting code for the given table data.
Figures D and  E display the prompts used to generate reasoning QA pairs for single-chart and multi-chart formats, respectively. We craft an example for each format to aid LLMs in understanding complex chart reasoning tasks and to generate step-by-step reasoning processes and precise answers that conform to the format. The executable plotting code is provided as auxiliary information to LLMs, making the generated QA pairs more reliable.
Figure F shows the prompt used for model evaluation. We employ GPT-4o to assess the match between the ground truth and the model’s predictions, where GPT-4o returns a score of 0 or 1 to indicate the correctness of the model’s prediction. Our evaluation focuses solely on the correctness of the final answer, disregarding the reasoning process.
Figure C: Prompt for code generation.
Figure D: Prompt for reasoning QA pairs generation for single-chart formats.
Figure E: Prompt for reasoning QA pairs generation for multi-chart formats.
Figure F: Prompt for ChartRQA evaluation using GPT-4o.